Tuesday, 18 August

Opening Ceremony

10:00–12:30 International Congress Center

W. A. Mozart, Clarinet Quintet K. 581, Movement 4

David Mumford, President of the International Mathematical Union Martin Grötschel, President of the ICM’98 Organizing Committee Karl-Heinz Hoffmann, President of the German Mathematical Society Friedrich Hirzebruch, Honorary President of the ICM’98 Organizing Committee

Paul Hindemith, Clarinet Quintet Op. 30, Movement 3–5

Wilhelm Staudacher, Head of the Office of the Federal President and State Secretary

Words of Welcome of the Federal President Roman Herzog Jürgen Rüttgers, Federal Minister of Education, Science, Research, and Technology Eberhard Diepgen, Governing Mayor of Berlin Hans-Jürgen Ewers, President of the Technical University Berlin Hansgeorg Hauser, Parliamentary State Secretary at the Federal Ministry of Finance

Presentation of a Commemorative Stamp on the Occasion of ICM’98 W. A. Mozart, Divertimento K. 287, Movement 6 Presentation of the Fields Medals and a Special Tribute

Yuri Manin, Chairman of the Fields Medal Committee Presentation of the Nevanlinna Prize David Mumford, Chairman of the Nevanlinna Prize Committee

(Music performed by soloists of the Ensemble Oriol Berlin)

Opening Reception

12:30–14:30 International Congress Center

This complimentary opening reception is offered to all participants of the Opening Ceremony. A light buffet lunch and beverages are served.

Laudationes

14:30–16:00 International Congress Center

Addresses on the work of the Fields medalists and the Nevanlinna Prize winner.

Plenary Lecture

16:15 – 17:15  International Congress Center

Jürgen Moser, ETH Zürich, Zürich, Switzerland

Dynamical Systems — Past and Present (I: p.17)

Chair: Friedrich Hirzebruch, Max-Planck-Institut für Mathematik, Bonn, Germany

ICM’98 – Program  1